the compound M(OH)x has a Ksp of 4*10^-6 and has a solubility of 10^-2 mol L^-1.the value of x should be

Grade:12

2 Answers

raj kumar
9 years ago
M(OH)x------M+x(OH) s s xs Ksp = (xs)^x/s 4*10^-6=(x*10^-2)^x/10^-2 4*10^-8= x^x*10^-2x thus x=4
Shruthi
5 years ago
Ksp=s(xs)^x.         M(OH)x------------>M+x(OH)
       4*10^-9=(10^-3(1+x))x^x
Option verifIcaton( when x= 2)
      4*10^-9=2^2(10^-3(1+2))
                    =4*10^-9